# Geometric Representation Condition Improves Equivariant Molecule Generation

### v1

@@ Abstract
Recent advancements in molecular generative models have demonstrated substantial potential in accelerating scientific discovery, particularly in drug design. However, these models often face challenges in generating high-quality molecules, especially in conditional scenarios where specific molecular properties must be satisfied. In this work, we introduce GeoRCG, a general framework to enhance the performance of molecular generative models by integrating geometric representation conditions. We decompose the molecule generation process into two stages: first, generating an informative geometric representation; second, generating a molecule conditioned on the representation. Compared to directly generating a molecule, the relatively easy-to-generate representation in the first-stage guides the second-stage generation to reach a high-quality molecule in a more goal-oriented and much faster way. Leveraging EDM as the base generator, we observe significant quality improvements in unconditional molecule generation on the widely-used QM9 and GEOM-DRUG datasets. More notably, in the challenging conditional molecular generation task, our framework achieves an average 31% performance improvement over state-of-the-art approaches, highlighting the superiority of conditioning on semantically rich geometric representations over conditioning on individual property values as in previous approaches. Furthermore, we show that, with such representation guidance, the number of diffusion steps can be reduced to as small as 100 while maintaining superior generation quality than that achieved with 1,000 steps, thereby significantly accelerating the generation process.
@@ 1 Introduction
Recent years have seen rapid development in generative modeling techniques for molecule generation (Garcia Satorras et al., 2021; Hoogeboom et al., 2022; Luo & Ji, 2022; Wu et al., 2022; Xu et al., 2023; Morehead & Cheng, 2024), which have demonstrated great promise in accelerating scientific discoveries such as drug design (Graves et al., 2020). By modeling molecules as point clouds of chemical elements embedded in Euclidean space and employing equivariant models as backbone architectures, such as EGNN (Satorras et al., 2021), these approaches can ensure the O(3)- (or SO(3)-) invariance of the modeled molecule probability and have shown significant promise in both unconditional and conditional molecule generations.
Despite the advances, precisely modeling the molecular distribution $q(\mathcal{M})$ still remains a challenge, with current models often falling short of satisfactory results. This is especially true in the more practical scenarios where the goal is to capture the conditional distribution $q(\mathcal{M}|c)$ for conditional generation, with $c$ representing a desired property such as the HOMO-LUMO gap. In such cases, recent models still produce molecules with property errors significantly larger than the data lower bound (Hoogeboom et al., 2022; Xu et al., 2023). Such challenge arises partly because molecules are naturally supported on a lower-dimensional manifold (Mislow, 2012; De Bortoli, 2022; You et al., 2023), yet are embedded in a 3D space with much higher ambient dimensions ($N\times(3+d)$, where $N$ is the number of atoms and $d$ the atom feature dimension). Consequently, directly learning these distributions without additional guidance or conditioning solely on a single property can result in substantial errors (Song et al., 2021), often leading to unstable or undesirable molecular samples.
Figure 1: Training and sampling procedure of GeoRCG in unconditional molecule generation scenario. a) During training, each molecule $\mathcal{M}$ is mapped into an informative representation $r$ by a pre-trained, frozen geometric encoder $E$. The distribution of representations is then learned by a lightweight representation generator. The molecule generator is trained in a self-conditioned manner, generating a molecule $\mathcal{M}$ conditioned on its own representation $E(\mathcal{M})$. b) During sampling, an informative representation is first generated, which subsequently guides the molecule generator to produce high-quality molecules.
@@
Our approach is directly inspired by RCG (Li et al., 2023), which, however, focuses on image data with fixed size and positions and does not necessitate handling Euclidean and permutation symmetries—factors that are markedly different in molecular data. Compared to recent work GraphRCG (Wang et al., 2024), which applies the RCG framework to 2D graph data, we explicitly handle 3D geometry that is more complex due to the additional Euclidean symmetry. Moreover, we avoid the complicated step-wise bootstrapped training and sampling process proposed in Wang et al. (2024) that requires noise alignment, sequential training, and simultaneous encoder training. Instead, we employ a simple and intuitive framework that enables parallel training and leverages advanced pre-trained geometric encoders containing valuable external knowledge (Zaidi et al., 2022; Feng et al., 2023), thus achieving competitive results while avoiding complex training procedures. Notably, while Li et al. (2023) primarily focus on empirical evaluation, we also provide theoretical characterizations of general representation-conditioned diffusion models for both unconditional and conditional generation, which offers a rigorous understanding of the improved performance.
To illustrate the effectiveness of our approach, we select one of the simplest and most classical equivariant generative models, EDM (Hoogeboom et al., 2022), as the base molecular generator of GeoRCG. Experimentally, our method achieves the following significant improvements:
  • Substantially enhancing the quality of the generated molecules on the widely used QM9 and GEOM-DRUG datasets. On QM9, GeoRCG not only improves the performance of EDM by a large margin, but also significantly surpasses several recent baselines with state-of-the-art performance (Wu et al., 2022; Xu et al., 2023; Morehead & Cheng, 2024; Song et al., 2024a).
  • More remarkably, in conditional molecule generation tasks, GeoRCG yields an average 31% improvement in performance, while many contemporary models struggled to achieve even marginal gains.
  • By incorporating classifier-free guidance in the molecule generator (Li et al., 2023) and employing low-temperature sampling for representation generation (Ingraham et al., 2023), GeoRCG demonstrate a flexible trade-off between molecular quality and diversity without additional training, which is especially advantageous in molecular generation tasks that prioritizes quality over diversity.
  • With the assistance of the representation guidance, GeoRCG can significantly reduce the number of diffusion steps needed, while preserving the molecular generation quality, thereby considerably accelerating the generation process.
@@ 2 Related Works
Early work has primarily focused on modeling molecules as 2D graphs (composed of atom types, connection, and edge types), utilizing 2D graph generative models to learn the graph distribution (Vignac et al., 2022; Jang et al., 2023; Jo et al., 2023; Luo et al., 2023; Zhou et al., 2024). However, since molecules inherently exist in 3D space, where physical laws govern their behavior and geometry provides critical information related to key properties, recent research has increasingly focused on utilizing 3D generative models to directly learn the geometric distribution by modeling molecules as point clouds of chemical elements. Notable early autoregressive models include G-SchNet (Gebauer et al., 2019) and G-SphereNet (Luo & Ji, 2022). More recently, diffusion models have demonstrated effectiveness in this domain, as evidenced by models like EDM (Hoogeboom et al., 2022) and subsequent advancements (Xu et al., 2023; Wu et al., 2022; Morehead & Cheng, 2024) that enhance EDM with latent space, prior information and more powerful backbones respectively. Furthermore, recent advances in flow methods (Lipman et al., 2022; Liu et al., 2022b) have inspired the development of geometric, equivariant flow methods including EquiFM (Song et al., 2024b) and GOAT (Hong et al., 2024), which can provide much faster molecule generation speed. Beyond these, there are also methods jointly model 2D and 3D information (Vignac et al., 2023; You et al., 2023; Irwin et al., 2024) (also called 3D graph (You et al., 2023)), where a representative method is MiDi (Vignac et al., 2023) which uses a diffusion framework to jointly diffuse and denoise atom type, bond type, formal charges and coordinates.
Latent Generative Models
@@
At a high level, our framework can also be viewed as a latent generative model, where data distributions are learned in a latent space (our stage 1) and decoded back through some decoder (our stage 2). Most prior works in this domain either focus on regular data forms with fixed positions and sizes (Van Den Oord et al., 2017; Razavi et al., 2019; Dai & Wipf, 2019; Aneja et al., 2021; Rombach et al., 2022; Li et al., 2023), or on data without Euclidean symmetry and require explicit modeling (Wang et al., 2024). Molecular data, however, presents unique challenges in both aspects. One of the key issues in this context is how to define the latent space—defining it as “latent coordinates and features” as in GeoLDM (Xu et al., 2023) still results in a geometrically structured and thus complex space, while defining it on representations as we do introduces the challenge of effectively “decoding” a global, non-symmetric embedding back into geometric objects. LGD (Zhou et al., 2024) trains a diffusion model on a unified Euclidean latent space obtained by jointly training a powerful encoder and a simple decoder, and performs both generation and prediction tasks focusing on 2D graphs. LDM-3DG (You et al., 2023) also adopts representation latent space but employs a cascaded (2D+3D) auto-encoder (AE) framework, where the decoder is designed (or trained) to be deterministic, rendering poor performance on the 3D part as evidenced in our experiments. In contrast, we model the decoder as a powerful generative model, focusing solely on geometric learning while demonstrating superior effectiveness. We leave the related works of pre-trained geometric encoders in Appendix A.
@@ 3.1 Preliminaries
In this work, we represent molecules as point clouds of chemical elements in 3D space, denoted by $\mathcal{M}=(\mathbf{x},\mathbf{h})$, where $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{R}^{N\times 3}$ represents the atomic coordinates of $N$ atoms, and $\mathbf{h}=(\mathbf{h}_{1},\dots,\mathbf{h}_{N})\in\mathbb{R}^{N\times d}$ captures the node features of dimension $d$, such as atomic numbers and charges. This formulation follows the approach of Hoogeboom et al. (2022); Xu et al. (2023); Morehead & Cheng (2024) and is widely utilized in molecular representation learning (Thomas et al., 2018; Li et al., 2024a; Zaidi et al., 2022), facilitating the integration of pre-trained molecular encoders (Zaidi et al., 2022; Feng et al., 2023). We use $q$ to denote the underlying data distribution, such as molecule distributions $q(\mathcal{M})$, and $p$ to denote the approximated distributions modeled by parametric methods.
@@ 3.2 GeoRCG: Geometric-Representation-Conditioned Molecular Generation
Geometric Representation Generator
To improve the quality of the generated molecules, we propose first transforming the geometrically structured molecular distribution $q(\mathcal{M})$ into a non-geometric representation distribution $q(r)$ using a well-pre-trained geometric encoder $E$ that maps each molecule $\mathcal{M}$ to its representation $r$. Learning the representation distribution $q(r)$ is considerably easier, since representations do not exhibit any symmetry as in explicit molecular generative models (Xu et al., 2022; Hoogeboom et al., 2022). We thus leverage a simple yet effective MLP-based diffusion architecture proposed in (Li et al., 2023) for the representation generator $p_{\varphi}(r)$, which adopts the DDIM architecture (Song et al., 2020) with MLP backbones and is optimized via the denoising score matching scheme (Vincent, 2011).
One deviation from previous practices (Li et al., 2023; Wang et al., 2024) is that we additionally condition the representation generator on the molecule’s node number $N$ by default [^1] We omit the condition $N$ in our probability decompositions and mathematical derivations for statement simplicity, as its inclusion does not affect the overall framework and conclusions.. This is crucial for ensuring consistency between the size of the representation’s underlying molecule and the size of the molecule it guides to generate. Moreover, molecules with different sizes often have distinct modes in structures and properties (Hoogeboom et al., 2022), which is reflected in their geometric representations learned by modern pre-trained geometric encoders (Zhou et al., 2023; Feng et al., 2023), as shown in Figure 2. From the figures, it is evident that by conditioning on $N$, the learning process for the representation generator becomes simpler and more effective, leading to the following loss function of our representation generator:
@@
There are several key properties of GeoRCG that facilitate high-quality molecule generation. First, GeoRCG preserves all symmetry properties of the base molecule generator $p_{\theta}(\mathcal{M})$:
@@
Moreover, representation-conditioned diffusion models can achieve no higher overall total variation distance than traditional diffusion models, and can arguably yield better results, as the representation encodes key data information that may further reduce estimation error. We present the rigorous bound in Theorem 3.1, and provide the proof and detailed discussions in Section E.1. Notably, this is a general theoretical characterization that applies to prior experimental work (Li et al., 2023).
@@
${\rm TV}(p_{\theta,\varphi}(x),q(x))\preceq$ $\underbrace{\sqrt{{\rm KL}(q_{0|\varphi}||\gamma^{d+3})}\exp(-T)}_{\text{convergence of forward process}}+\underbrace{(L_{x}\sqrt{(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}$ (3)
@@
Balancing Quality and Diversity of Molecule Generation
In many scientific applications, researchers prioritize generating higher-quality molecules over more diverse ones. To address this preference, we introduce a feature that allows fine-grained control over the trade-off between diversity and quality in the sampling stage (thus without retraining). This is achieved by integrating two key techniques: low-temperature sampling (Ho & Salimans, 2022) (controlled via the temperature $\mathcal{T}$) for the representation generator, and classifier-free guidance (Ho & Salimans, 2022) (controlled via the coefficient $w$) for the molecule generator. We provide more details about the two techniques in Appendix B. Combining the two features enables a flexible and explicit control, which we refer to as “Balancing Controllablility” and demonstrate its effectiveness in Section 4.2.
Handling Conditional Molecule Generation
@@
4 Experiments
4.1 Experiment Setup
Datasets and Tasks
As a method for 3D molecule generation, we evaluate GeoRCG on the widely used datasets QM9 (Ramakrishnan et al., 2014) and GEOM-DRUG (Gebauer et al., 2019; 2022). We focus on two tasks: unconditional molecule generation, where the goal is to sample from $q(\mathcal{M})$, and conditional (or controllable) molecule generation, where a property $c$ is given, and we aim to sample from $q(\mathcal{M}|c)$. To ensure fair comparisons, we follow the dataset split and configurations exactly as in Anderson et al. (2019); Hoogeboom et al. (2022); Xu et al. (2023). Without further clarification, we bold the highest scores and underline the second-highest one. Additionally, to illustrate the direct improvement over our base model, EDM (Hoogeboom et al., 2022), we display green numbers next to the score to indicate the average improvement, and red numbers to denote a decrease. All results are calculated based on 10k randomly sampled molecules, averaged over three runs, with standard errors reported in parentheses.
Instantiation of the Pre-trained Encoder
We employ Frad (Feng et al., 2023), which was pre-trained on the PCQM4Mv2 dataset (Nakata & Shimazaki, 2017) using a hybrid noise denoising objective, as the geometric encoder for QM9 dataset. For GEOM-DRUG, we adopt Unimol (Zhou et al., 2023) and pre-train it on the GEOM-DRUG dataset itself, since GEOM-DRUG contains distinct chemical elements not present in PCQM4Mv2 or other commonly used pre-training datasets such as ZINC or ChemBL (Li et al., 2021).
Baselines
A direct comparison is made with our base molecule generator, EDM (Hoogeboom et al., 2022). Additionally, we benchmark against the non-equivariant counterparts of EDM and GeoLDM (Xu et al., 2023), namely GDM(-AUG) (Hoogeboom et al., 2022) and GraphLDM(-AUG) (Xu et al., 2023), as well as the autoregressive method G-SchNet (Gebauer et al., 2019). For further comparison, we include advanced equivariant diffusion models like GeoLDM (Xu et al., 2023), EDM-Bridge (Wu et al., 2022), and GCDM (Morehead & Cheng, 2024), along with fast equivariant flow-based methods such as E-NF (Garcia Satorras et al., 2021), EquiFM (Song et al., 2024b), and GOAT (Hong et al., 2024). A recent Bayesian-based method GeoBFN (Song et al., 2024a) is also considered for its high-quality samples. Finally, we extend the comparison to 2D and 3D methods like MiDi (Vignac et al., 2023) and LDM-3DG (You et al., 2023).
4.2 Unconditional Molecule Generation

### v2

@@ Abstract
Recent advancements in molecular generative models have demonstrated substantial potential in accelerating scientific discovery, particularly in drug design. However, these models often face challenges in generating high-quality molecules, especially in conditional scenarios where specific molecular properties must be satisfied. In this work, we introduce GeoRCG, a general framework to enhance the performance of molecular generative models by integrating geometric representation conditions with provable theoretical guarantees. We decompose the molecule generation process into two stages: first, generating an informative geometric representation; second, generating a molecule conditioned on the representation. Compared to directly generating a molecule, the relatively easy-to-generate representation in the first stage guides the second-stage generation to reach a high-quality molecule in a more goal-oriented and much faster way. Leveraging EDM (Hoogeboom et al., 2022) and SemlaFlow (Irwin et al., 2024) as the base generators, we observe significant quality improvements in unconditional molecule generation tasks on the widely-used QM9 and GEOM-DRUG datasets. More notably, in the challenging conditional molecular generation task, our framework achieves an average 31% performance improvement over state-of-the-art approaches, highlighting the superiority of conditioning on semantically rich geometric representations over conditioning on individual property values as in previous approaches. Furthermore, we show that, with such representation guidance, the number of diffusion steps can be reduced to as small as 100 while largely preserving the generation quality achieved with 1,000 steps, thereby significantly accelerating the generation process.
Machine Learning, ICML
@@ 1 Introduction
Recent years have seen rapid development in generative modeling techniques for molecule generation (Garcia Satorras et al., 2021; Hoogeboom et al., 2022; Luo & Ji, 2022; Wu et al., 2022; Xu et al., 2023; Le et al., 2023; Morehead & Cheng, 2024), which have demonstrated great promise in accelerating scientific discoveries such as drug design (Graves et al., 2020). By representing molecules as point clouds of chemical elements embedded in Euclidean space (potentially with edges (Vignac et al., 2023; Irwin et al., 2024)) and employing equivariant models as backbone architectures, such as EGNN (Satorras et al., 2021), these approaches can ensure the O(3)- (or SO(3)-) invariance of the modeled molecule probability and have shown significant promise in both unconditional and conditional molecule generations.
Despite the advances, precisely modeling the molecular distribution $q(\mathcal{M})$ still remains a challenge, with current models often falling short of satisfactory results. This is especially true in the more practical scenarios where the goal is to capture the conditional distribution $q(\mathcal{M}|c)$ for conditional generation, with $c$ representing a desired property such as the HOMO-LUMO gap. In such cases, recent models still produce molecules with property errors significantly larger than the data lower bound (Hoogeboom et al., 2022; Xu et al., 2023). This challenge arises in part because molecules are naturally supported in a lower-dimensional manifold (Mislow, 2012; De Bortoli, 2022; You et al., 2023), yet they are embedded in a 3D space with much higher ambient dimensions ($N\times(3+d)$, where $N$ is the number of atoms and $d$ the atom feature dimension). Consequently, directly learning these distributions without additional guidance or conditioning solely on a single property can result in substantial errors (Song et al., 2021), often leading to unstable or undesirable molecular samples.
Figure 1: Training and sampling procedure of GeoRCG for unconditional molecule generation. a) During training, each molecule $\mathcal{M}$ is mapped into a representation $r$ by a pre-trained, frozen geometric encoder $E$. The representation distribution is then learned by a lightweight representation generator. The molecule generator is trained in a self-conditioned manner, generating a molecule $\mathcal{M}$ conditioned on its own representation $E(\mathcal{M})$. b) During sampling, an informative representation is first generated, which subsequently guides the molecule generator to produce high-quality molecules.
@@
Our approach is directly inspired by RCG (Li et al., 2023), which, however, focuses on image data with fixed size and positions and does not necessitate handling Euclidean and permutation symmetries—factors that are markedly different in molecular data. Compared to recent work GraphRCG (Wang et al., 2024), which applies the RCG framework to 2D graph data, we explicitly handle 3D geometry that is more complex due to the additional Euclidean symmetry. Moreover, we avoid the complicated step-wise bootstrapped training and sampling process proposed in Wang et al. (2024) that requires noise alignment, sequential training, and simultaneous encoder training. Instead, we employ a simple and intuitive framework that enables parallel training and leverages advanced pre-trained geometric encoders containing valuable external knowledge (Zaidi et al., 2022; Feng et al., 2023), thus achieving competitive results while avoiding complex training procedures. Notably, while Li et al. (2023) primarily focus on empirical evaluation, we also provide generic theoretical characterizations of the representation-conditioned diffusion model class for both unconditional and conditional generation, which offers a rigorous understanding of the improved performance.
To illustrate the effectiveness of our approach, we select one of the simplest and most classical equivariant generative models, EDM (Hoogeboom et al., 2022), as the base molecular generator of GeoRCG. For better performance on the more challenging dataset GEOM-DRUG, we also adopt GeoRCG onto recent state-of-the-art (SOTA) model SemlaFlow (Irwin et al., 2024). Experimentally, our method achieves the following significant improvements:
  • Substantially enhancing the quality (e.g., molecule stability) of the generated molecules on the widely used QM9 and GEOM-DRUG datasets. On QM9, GeoRCG not only improves the performance of EDM by a large margin, but also significantly surpasses several recent baselines with advanced performance (Wu et al., 2022; Xu et al., 2023; Morehead & Cheng, 2024; Song et al., 2024a). On GEOM-DRUG, GeoRCG consistently enhances SemlaFlow’s already SOTA performance.
  • More remarkably, in conditional molecule generation tasks, GeoRCG yields an average 31% improvement in performance (i.e., difference of generated molecule’s property with conditions), while many contemporary models struggled to achieve even marginal gains.
  • By incorporating classifier-free guidance into the molecule generator (Li et al., 2023) and employing low-temperature sampling for representation generation (Ingraham et al., 2023), GeoRCG demonstrate a flexible trade-off between molecular quality and diversity on QM9 dataset without additional training, which is especially advantageous in molecular generation tasks that prioritize quality over diversity.
  • With the assistance of the representation guidance, GeoRCG can significantly reduce the number of diffusion steps required by approximately 10x, while preserving the quality of molecular generation, thus considerably accelerating the generation process.
@@ 2 Related Works
Molecular Generative Models. Early work has primarily focused on modeling molecules as 2D graphs (composed of atom types, connection, and edge types), utilizing 2D graph generative models to learn the graph distribution (Vignac et al., 2022; Jang et al., 2023; Le et al., 2023; Jo et al., 2023; Luo et al., 2023; Zhou et al., 2024). However, since molecules inherently exist in 3D space, where physical laws govern their behavior and geometry provides critical information related to key properties, recent research has increasingly focused on utilizing 3D generative models to directly learn the geometric distribution by modeling molecules as point clouds of chemical elements. Notable early autoregressive models include G-SchNet (Gebauer et al., 2019) and G-SphereNet (Luo & Ji, 2022). More recently, diffusion models have demonstrated effectiveness in this domain, as evidenced by models like EDM (Hoogeboom et al., 2022) and subsequent advancements (Xu et al., 2023; Wu et al., 2022; Morehead & Cheng, 2024) that enhance EDM with latent space, prior information and more powerful backbones respectively. Furthermore, recent advances in flow methods (Lipman et al., 2022; Liu et al., 2022b) have inspired the development of geometric, equivariant flow methods including EquiFM (Song et al., 2024b) and GOAT (Hong et al., 2024), which can provide much faster molecule generation speed. Beyond these, there are also methods jointly model 2D and 3D information (Vignac et al., 2023; You et al., 2023; Huang et al., 2024; Irwin et al., 2024) (also called 3D graph (You et al., 2023)), where a representative method is MiDi (Vignac et al., 2023) which uses a diffusion framework to jointly diffuse and denoise atom type, bond type, formal charges and coordinates.
@@
Latent Generative Models. At a high level, our framework can also be viewed as a latent generative model, where data distributions are learned in a latent space (our stage 1) and decoded back through some decoder (our stage 2). Most prior work in this domain either focuses on regular data forms with fixed positions and sizes (Van Den Oord et al., 2017; Razavi et al., 2019; Dai & Wipf, 2019; Aneja et al., 2021; Rombach et al., 2022; Li et al., 2023), or on data without Euclidean symmetry and requires explicit modeling (Wang et al., 2024). Molecular data, however, presents unique challenges in both aspects. One of the key issues in this context is how to define the latent space—defining it as “latent coordinates and features” as in GeoLDM (Xu et al., 2023) still results in a geometrically structured and thus complex space, while defining it on representations as we do introduces the challenge of effectively “decoding” a global, non-symmetric embedding back into geometric objects. LGD (Zhou et al., 2024) trains a diffusion model on a unified Euclidean latent space obtained by jointly training a powerful encoder and a simple decoder, and performs both generation and prediction tasks focusing on 2D graphs. LDM-3DG (You et al., 2023) also adopts representation latent space but employs a cascaded (2D+3D) auto-encoder (AE) framework, where the decoder is designed (or trained) to be deterministic, rendering poor performance on the 3D part as evidenced in our experiments. In contrast, we model the decoder as a powerful generative model, focusing solely on geometric learning while demonstrating superior effectiveness.
We leave the related works of pre-trained geometric encoders in Appendix A.
@@ 3.1 Preliminaries
In this work, we represent molecules as point clouds of chemical elements in 3D space, denoted by $\mathcal{M}=(\mathbf{x},\mathbf{h})$, where $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{R}^{N\times 3}$ represents the atomic coordinates of $N$ atoms, and $\mathbf{h}=(\mathbf{h}_{1},\dots,\mathbf{h}_{N})\in\mathbb{R}^{N\times d}$ captures the node features of dimension $d$, such as atomic numbers and charges. This formulation follows the approach of Hoogeboom et al. (2022); Xu et al. (2023); Morehead & Cheng (2024) and is widely utilized in molecular representation learning (Thomas et al., 2018; Li et al., 2024a; Zaidi et al., 2022), facilitating the integration of pre-trained molecular encoders (Zaidi et al., 2022; Feng et al., 2023). After generating point clouds of chemical elements, these methods infer bond types using lookup tables based on atom types and pairwise distances, or relying on advanced packages like OpenBabel (O’Boyle et al., 2011). Notably, approaches like MiDi (Vignac et al., 2023) and SemlaFlow (Irwin et al., 2024) additionally represent molecules with explicit bond types, enabling joint learning and generating of 2D and 3D information, which typically results in improved performance. We use $q$ to denote the underlying data distribution, such as molecule distributions $q(\mathcal{M})$, and $p$ to denote the approximated distributions modeled by parametric methods.
@@ 3.2 GeoRCG: Geometric-Representation-Conditioned Molecular Generation
Geometric Representation Generator. To improve the quality of the generated molecules, we propose first transforming the geometrically structured molecular distribution $q(\mathcal{M})$ into a non-geometric representation distribution $q(r)$ using a well-pre-trained geometric encoder $E$ that maps each molecule $\mathcal{M}$ to its representation $r$. Learning the representation distribution $q(r)$ is considerably easier, since representations do not exhibit any symmetry as in explicit molecular generative models (Hoogeboom et al., 2022). We thus leverage a simple yet effective MLP-based diffusion architecture proposed in (Li et al., 2023) for the representation generator $p_{\varphi}(r)$, which adopts the DDIM architecture (Song et al., 2020) with MLP backbones and is optimized via the denoising score matching scheme (Vincent, 2011).
One deviation from previous practices (Li et al., 2023; Wang et al., 2024) is that we additionally condition the representation generator on the molecule’s node number $N$ by default [^1] We omit the condition $N$ in our probability decompositions and mathematical derivations for statement simplicity, as its inclusion does not affect the overall framework and conclusions.. This is crucial for ensuring consistency between the size of the representation’s underlying molecule and the size of the molecule it guides to generate. Moreover, molecules with different sizes often have distinct modes in structures and properties (Hoogeboom et al., 2022), which is reflected in their geometric representations learned by modern pre-trained geometric encoders (Zhou et al., 2023; Feng et al., 2023), as shown in Figure 4 in Appendix D. From the figures, it is evident that by conditioning on $N$, the learning process for the representation generator becomes simpler and more effective, leading to the following loss function of our representation generator:
@@
EDM is designed to ensure the $O(3)$-invariance, i.e., for any $\mathbf{R}\in O(3)$, $p_{\theta}(\mathcal{M})=p_{\theta}(\mathbf{x},\mathbf{h}))=p_{\theta}((\mathbf{x}\mathbf{R}^{T},\mathbf{h}))$. To accomodate EDM to representation conditions, we use the following training objective:
@@
There are several key properties of GeoRCG that facilitate high-quality molecule generation. First, GeoRCG preserves symmetry properties of the base molecule generator $p_{\theta}(\mathcal{M})$:
@@
Moreover, representation-conditioned diffusion models can achieve no higher overall total variation distance than traditional diffusion models, and can arguably yield better results, as the representation encodes key data information that may further reduce estimation error. We present the rigorous bound in Theorem 3.2, and provide the proof and detailed discussions in Section E.1. Notably, this is a generic theoretical characterization that applies to prior experimental work (Li et al., 2023). For models that account for equivariant symmetries, such as EDM, we build upon results from (Feng et al., 2024; You et al., 2023) to establish finer-grained bounds, as detailed in Theorem E.14.
@@
${\rm TV}(p_{\theta,\varphi}(x),q(x))\preceq$ $\underbrace{\sqrt{{\rm KL}(q_{0|\varphi}||\gamma^{N(d+3)})}\exp(-T)}_{\text{convergence of forward process}}$ (3)
$+\underbrace{(L_{x}\sqrt{N(d+3)h}+L_{x}m_{x}h)\sqrt{T}}_{\text{discretization error}}$ (4)
$+\underbrace{\epsilon_{\varphi,\theta,\rm cond}\sqrt{T}}_{\text{conditional score estimation error}}$ (5)
$+\underbrace{{\rm TV}(q_{0|\varphi},q_{0})}_{\text{representationgeneration error}}$ (6)
Balancing Quality and Diversity of Molecule Generation. In many scientific applications, researchers prioritize generating higher-quality molecules over more diverse ones. To facilitate this, we introduce a feature that allows fine-grained control over the trade-off between diversity and quality in the sampling stage (thus without retraining). This is achieved by integrating two key techniques: low-temperature sampling (Ingraham et al., 2023) (controlled via the temperature $\mathcal{T}$) for the representation generator, and classifier-free guidance (Ho & Salimans, 2022; Zheng et al., 2023) (controlled via the coefficient $w$) for the molecule generator. We provide more details about the two techniques in Appendix B. Combining the two features enables a flexible and explicit control, which we refer to as “Balancing Controllablility” and demonstrate its effectiveness in Section 4.2.
@@
4 Experiments
4.1 Experiment Setup
Datasets and Tasks. As a method for 3D molecule generation, we evaluate GeoRCG on the widely used datasets QM9 (Ramakrishnan et al., 2014) and GEOM-DRUG (Gebauer et al., 2019, 2022; Axelrod & Gomez-Bombarelli, 2022). We focus on two tasks: unconditional molecule generation, where the goal is to sample from $q(\mathcal{M})$, and conditional (or controllable) molecule generation, where a property $c$ is given, and we aim to sample from $q(\mathcal{M}|c)$. By default, “GeoRCG” refers to our method with EDM (Hoogeboom et al., 2022) as base model, while its application to SemlaFlow is explicitly labeled as “GeoRCG (Semla)”. To ensure fair comparisons, we follow the dataset split and configurations exactly as in Anderson et al. (2019); Hoogeboom et al. (2022); Xu et al. (2023); Vignac et al. (2023). Without further clarification, we bold the highest scores and underline the second-highest one. Additionally, to illustrate the direct improvement over the base model, we display green numbers next to the score to indicate the average improvement, and red numbers to denote a decrease. Without further clarification, results are calculated based on 10k randomly sampled molecules, averaged over three runs, with standard errors reported in parentheses.
Instantiation of the Pre-trained Encoder. We employ Frad (Feng et al., 2023), which was pre-trained on the PCQM4Mv2 dataset (Nakata & Shimazaki, 2017) using a hybrid noise denoising objective, as the geometric encoder for QM9 dataset. For GEOM-DRUG, we adopt Unimol (Zhou et al., 2023) and additionally add GEOM-DRUG to its pre-training dataset, since GEOM-DRUG contains distinct chemical elements not present in PCQM4Mv2 or other commonly used pre-training datasets such as ZINC or ChemBL (Li et al., 2021).
Baselines. A direct comparison is made with our base molecule generators, EDM (Hoogeboom et al., 2022) or SemlaFlow (Irwin et al., 2024). For GeoRCG applied to EDM, we compare it against generative models that, like EDM, do not explicitly generate bonds but instead infer them based on bond lengths. Although this approach may be less effective in generating valid molecules, it is widely adopted and presents a greater challenge for generative models in learning 3D geometric distributions—precisely where our geometric representation guidance offers the most significant improvement. These models include: (1) the non-equivariant counterparts of EDM and GeoLDM (Xu et al., 2023), specifically GDM(-AUG)(Hoogeboom et al., 2022) and GraphLDM(-AUG)(Xu et al., 2023); (2) the autoregressive method G-SchNet (Gebauer et al., 2019); (3) advanced equivariant diffusion models such as GeoLDM (Xu et al., 2023), EDM-Bridge (Wu et al., 2022), and GCDM (Morehead & Cheng, 2024); (4) fast equivariant flow-based methods like E-NF (Garcia Satorras et al., 2021), EquiFM (Song et al., 2024b), and GOAT (Hong et al., 2024); and (5) the recently introduced Bayesian-based method GeoBFN (Song et al., 2024a). For GeoRCG (Semla), we compare it with recent advanced 2D&3D methods that directly generate bonds to produce higher-quality samples similar to SemlaFlow, including MiDi (Vignac et al., 2023) and EQGAT-diff (Le et al., 2023).
4.2 Unconditional Molecule Generation